OpenStudy (anonymous):

Determine the geometric partial sum is convergent or divergent? can you show me the steps?

OpenStudy (anonymous):

Hmm I don't know how to do that actually? If you could post a problem, maybe then I'll be able to help you

OpenStudy (anonymous):

lol sorry about tht.

OpenStudy (anonymous):

any finite sum is a number so it "converges"

OpenStudy (anonymous):

if the question is does \[\sum_{k=1}^{\infty} 3(5^k)=\lim_{n\rightarrow \infty}\sum_{k=1}^n 3\times 5^k\] converge the answer is surely not because the terms do not even go to zero, they get bigger and bigger!